OpenStudy (anonymous):

Select the answer choice below that represents 5.9 x 10-4 in standard notation. 0.000059 0.00059 59,000 5,900

OpenStudy (anonymous):

what is \[10^{-4}\]?

mathslover (mathslover):

\[5.9 * 10^{-4}\] this is written in scientific notation , to write it in standard notation it will be written as \[59/10 * 1/10^4 = 59/10^5 = 0.00059 \]

mathslover (mathslover):

hence the correct answer is b